# Towards Anytime Optical Flow Estimation with Event Cameras

### v1

@@ IV-A Datasets
DSEC. The DSEC dataset, introduced by Gehrig et al. [12], consists of 24 sequences captured in real-world outdoor driving scenes. The dataset includes a total of 7,800 training samples and 2,100 testing samples, captured at a resolution of $640{\times}480$. It covers both daytime and nighttime scenarios, encompassing small and large optical flows. Moreover, the DSEC dataset offers high-precision sparse ground truth values for optical flow. As the dataset does not include an official validation set, we divided the training dataset randomly using a fixed seed. To create the training and validation sets, we allocated them in a $4{:}1$ ratio. Consequently, only $80\%$ of the training data was utilized for training our model.
@@
MVSEC. MVSEC [14] is a classic real event optical flow dataset that comprises sequences from various indoor and outdoor scenes with a resolution of $346{\times}260$. The event density of the dataset is relatively sparse, and the optical flow ground truth distribution mostly highlights small flows. As a result, two different time intervals, i.e., $dt{=}1$ and $dt{=}4$, are used for accuracy evaluation. $dt{=}1$ means that adjacent grayscale frames are used as a sample, whereas $dt{=}4$ represents a four-fold increase in the grayscale frame interval. We follow the E-RAFT convention [11], conduct training solely on the outdoor day 2 sequence, and evaluate on $800$ samples data from outdoor day 1.
@@ IV-B Implementation Details
For the DSEC dataset [12], we use the Adam optimizer with a batch size of $6$ and a learning rate of $5{\times}1e^{-4}$ to train for $100k$ iterations on the training set we divided, which accounts for $80\%$ of the total dataset. Then, we reduce the learning rate by a factor of $10$ and continue training for another $10k$ iterations. We applied the same training settings to networks employing different numbers of bins. During the model training on the DSEC dataset, two online data augmentation methods, namely random cropping and horizontal flipping, were employed. Random cropping was performed with a size of $288{\times}384$, while horizontal flipping was applied with a probability of $50\%$.
@@
For the MVSEC dataset [14], we employ the Adam optimizer with a batch size of $3$ and a learning rate of $1{\times}1e^{-3}$ to train for $120k$ iterations. We utilize random cropping with a size of $256{\times}256$ and horizontal flipping with a probability of $50\%$ during training. For MVSEC scenes with $dt{=}1$, we use the setting with $bins{=}3$; for MVSEC with $dt{=}4$, we use the setting with $bins{=}9$.
TABLE 1: Evaluation results on the DSEC hidden test set [12].
Methods EPEnormal-$\downarrow$ AEnormal-$\downarrow$ 1PEnormal-$\downarrow$ 3PEnormal-$\downarrow$ Runtime
(ms/frame)
EV-FlowNet [14] 2.32 7.90 55.4 18.6 18
E-RAFT [11] 0.82 - 13.5 3.0 93
IDNet 1 iter [39] 1.30 4.82 33.7 6.7 -
TIDNet† [39] 0.93 3.98 19.3 3.4 -
EVA-Flow (Ours) 0.88 3.31 15.9 3.2 9.2
E-RAFT† [11] 0.79 2.85 12.7 2.7 93
† denotes that the method uses the warm-start strategy.
IV-C Regular Flow Evaluation Prototype
Evaluation on DSEC. Tab. 1 displays the evaluation results on the DSEC flow benchmark [12]. It is worth noting that this approach is the only optical flow estimation method that has ultra-low data latency ($5ms$) and a high output frame rate ($200Hz$), while also achieving competitive accuracy compared to state-of-the-art methods. This approach can estimate the current optical flow as soon as the event data arrives every $5ms$, without needing to wait for the entire sample period ($100ms$) to be completed, resulting in extremely low data latency and a super high frame rate($200Hz$). In contrast, other approaches, constrained by the frame rate of the ground truth dataset, can only provide optical flow outputs at 10 Hz. Furthermore, the computational cost of a single optical flow estimation in this approach is remarkably low. Our runtime for a single optical flow estimation is the lowest among other methods, with only $10\%$ of E-RAFT’s runtime. (The runtime was obtained by testing on the 2080Ti.)
@@ IV-C Regular Flow Evaluation Prototype
The qualitative results of our method on the DSEC flow dataset are shown in Fig. 6. To make a more intuitive comparison of the accuracy of optical flow estimation, we also visualized the motion-compensated (MC) event frames. The sharper the MC frames, the more precise the optical flow estimates. It can be seen from the Regions of Interest (ROIs) in Fig. 6 that, compared with E-RAFT, our model can better distinguish the outline of small objects (such as poles) and obtain more accurate optical flow estimates. Our analysis indicates that the implicit use of warp-aligned edges of events in our approach to obtain motion features is the distinguishing factor compared to E-RAFT, which instead uses correlation volume. For small objects with simple and clear contours, it is relatively easy to warp and align event features. However, due to the limited textures, the correlation-based method, which relies on measuring feature similarity to obtain motion features, fails to obtain high-quality motion features when there are no sufficient texture features, resulting in the inferior performance of E-RAFT in these areas.
TABLE 2: Evaluation results on MVSEC [14].
$dt{=}1$ EPE $\downarrow$ Outlier%$\downarrow$
SSL EV-FlowNet [14] 0.49 0.20
Spike-FlowNet [47] 0.49 -
STE-FlowNet [48] 0.42 0.00
USL Hagenaars et al. [49] 0.47 0.25
Zhu et al. [17] 0.32 0.00
SL E-RAFT [11] 0.24 0.00
EVA-Flow (Ours) 0.25 0.00
E-RAFT [11] (Zero-Shot) 0.53 1.42
EVA-Flow (Zero-Shot) 0.39 0.07
$dt{=}4$
SSL EV-FlowNet [14] 1.23 7.30
Spike-FlowNet [47] 1.09 -
STE-FlowNet [48] 0.99 3.90
USL Hagenaars et al. [49] 1.69 12.5
Zhu et al. [17] 1.30 9.70
SL E-RAFT [11] 0.72 1.12
EVA-Flow (Ours) 0.82 2.41
E-RAFT [11] (Zero-Shot) 1.93 17.7
EVA-Flow (Zero-Shot) 0.96 4.92
TABLE 3: Comparison of anytime flow estimation on DSEC [12].
Sequence Method RFWL Avg. Runtime (ms/frame)
$\mathbf{V}_{0\xrightarrow{}1}$ $\mathbf{V}_{0\xrightarrow{}2}$ $\mathbf{V}_{0\xrightarrow{}3}$ $\mathbf{V}_{0\xrightarrow{}4}$ $\mathbf{V}_{0\xrightarrow{}5}$ $\mathbf{V}_{0\xrightarrow{}6}$ $\mathbf{V}_{0\xrightarrow{}7}$ $\mathbf{V}_{0\xrightarrow{}8}$ $\mathbf{V}_{0\xrightarrow{}9}$ $\mathbf{V}_{0\xrightarrow{}10}$ $\mathbf{V}_{0\xrightarrow{}11}$ $\mathbf{V}_{0\xrightarrow{}12}$ $\mathbf{V}_{0\xrightarrow{}13}$ $\mathbf{V}_{0\xrightarrow{}14}$
thun_01_a E-RAFT* 1.039 1.104 1.163 1.211 1.250 1.282 1.309 1.333 1.355 1.374 1.392 1.407 1.422 1.434 1.291 93
EVA-Flow (Ours) 1.030 1.101 1.166 1.217 1.256 1.289 1.316 1.339 1.359 1.379 1.394 1.408 1.421 1.434 1.294 9.2
thun_01_b E-RAFT* 1.058 1.134 1.193 1.236 1.277 1.313 1.349 1.382 1.416 1.447 1.476 1.500 1.522 1.539 1.346 93
EVA-Flow (Ours) 1.046 1.135 1.208 1.262 1.305 1.340 1.370 1.393 1.419 1.444 1.467 1.490 1.512 1.532 1.352 9.2
interlaken_01_a E-RAFT* 1.098 1.218 1.316 1.406 1.483 1.552 1.624 1.694 1.758 1.816 1.868 1.911 1.947 1.970 1.619 93
EVA-Flow (Ours) 1.084 1.218 1.332 1.437 1.519 1.585 1.645 1.705 1.758 1.810 1.857 1.898 1.934 1.961 1.625 9.2
interlaken_00_b E-RAFT* 1.105 1.233 1.338 1.426 1.500 1.565 1.621 1.670 1.714 1.753 1.788 1.819 1.843 1.860 1.588 93
EVA-Flow (Ours) 1.088 1.229 1.340 1.437 1.515 1.576 1.626 1.670 1.709 1.742 1.770 1.796 1.816 1.837 1.582 9.2
zurich_city_12_a E-RAFT* 1.005 1.020 1.034 1.050 1.065 1.079 1.090 1.103 1.116 1.127 1.139 1.149 1.161 1.169 1.093 93
EVA-Flow (Ours) 1.004 1.021 1.032 1.046 1.060 1.075 1.085 1.099 1.111 1.122 1.134 1.145 1.157 1.166 1.090 9.2
zurich_city_14_c E-RAFT* 1.057 1.155 1.250 1.329 1.391 1.453 1.510 1.555 1.598 1.636 1.666 1.695 1.724 1.752 1.484 93
EVA-Flow (Ours) 1.044 1.152 1.249 1.333 1.394 1.460 1.517 1.561 1.605 1.645 1.675 1.704 1.732 1.761 1.488 9.2
zurich_city_15_a E-RAFT* 1.071 1.174 1.256 1.324 1.384 1.436 1.480 1.522 1.566 1.606 1.642 1.676 1.703 1.721 1.469 93
EVA-Flow (Ours) 1.062 1.177 1.270 1.345 1.409 1.461 1.502 1.542 1.580 1.614 1.646 1.677 1.701 1.720 1.479 9.2
E-RAFT* indicates that the intermediate flow is obtained by interpolating the E-RAFT predictions over the complete time range.
@@
Figure 6: Qualitative examples of optical flow predictions on the DSEC test sequences [12]. The four samples are obtained from different scenes belonging to distinct sequences. The areas of interest are demarcated by red boxes in the images. These regions are enlarged by a factor of two and presented in the upper left corner of the corresponding images. The abbreviation MC represents motion-compensated.
@@
Figure 7: Qualitative examples of optical flow predictions on the outdoor day 1 sequence of the MVSEC dataset [14] ($dt{=}4$). The model used here is only trained on the DSEC dataset [12]. The respective EPE and RFWL metrics can be found in the bottom left corner of the images. The areas of interest are denoted as red boxes in the images.
@@
Evaluation on the MVSEC dataset. The evaluation results of the MVSEC benchmark [14] are presented in Tab. 2. SSL, USL, and SL stand for self-supervised-, unsupervised-, and supervised learning, respectively. While our model’s accuracy is slightly lower than that of E-RAFT based on the data in the table, it is worth highlighting that our zero-shot results display a significantly better accuracy when directly testing the performance of the DSEC-trained model on the MVSEC dataset. Especially noteworthy is our model’s performance at $dt{=}4$, where our model’s EPE is halved compared to that of E-RAFT ($0.96$ vs. $1.93$). These results demonstrate the superior generalizability of the proposed EVA-Flow.
@@
Figure 8: Qualitative examples of optical flow predictions on our EVA-FlowSet. The model used here is only trained on the DSEC dataset [12]. The initial three rows of samples are obtained from high-speed sequences, while the remaining three rows are taken from normal-speed ones.
Figure 9: Evaluation results of time-dense optical flow using the RFWL in EVA-FlowSet. Models utilized here are exclusively trained on the DSEC dataset [12]. Our proposed model significantly outperforms E-RAFT in fast-motion scenarios.
@@
The qualitative results of testing MVSEC under the zero-shot setting, i.e., the model is only trained on DSEC, can be found in Fig. 7. Owing to the small optical flow of MVSEC, the motion compensation effect is indistinct under the $dt{=}1$ setting. Hence, we apply a longer time range of grayscale frames, i.e., $dt{=}4$. There is a marked discrepancy in motion speed between the MVSEC and DSEC datasets, as depicted in Fig. 7 and Fig. 6. More specifically, the event images in the MVSEC dataset display minimal motion blur caused by slow movement, whereas the DSEC dataset showcases significant motion blur attributed to rapid motion. Nonetheless, our model is capable of achieving near-ground-truth optical flow results on the MVSEC dataset, while E-RAFT shows inadequate generalization performances, particularly in non-event ground areas. Additionally, our model demonstrates superior capabilities for small targets such as poles and road markings, accurately depicting their contours with improved motion compensation results.
@@
Evaluation on EVA-FlowSet. Qualitative results of the curve motion scene dataset EVA-FlowSet can be found in Fig. 8. By visualizing the optical flow images, it can be observed that, for both fast-moving scenes and normal scenes, our model can estimate the contour of moving objects more accurately than E-RAFT [11]. The magnified ROI area shows that our model obtains sharper motion compensation results in the first three rows of fast-moving scenes, while E-RAFT’s motion compensation performance is comparable to our model in the last three rows of moderate-moving-speed scenes. E-RAFT calculates the correlation between the event voxel grid before and after the start time to estimate the optical flow, which implicitly assumes that the optical flow is constant throughout the entire time. In contrast, our method uses the SMR module iteratively in each time step to obtain continuous optical flow estimation, which is suitable for curve motion scenes. For moderate-moving-speed scenes, even if the motion trajectory is curved, the motion could be approximated as linear due to the slow motion speed. In moderate-moving-speed cases, E-RAFT’s motion compensation result is comparable to our model. However, in fast-moving scenes, the linear motion assumption of E-RAFT is no longer valid due to the curved nature of the motion trajectory. This scenario is particularly critical for safety concerns. In contrast, our model utilizes a dense optical flow estimation framework that does not depend on the linear motion assumption. Consequently, our model significantly outperforms E-RAFT in fast-moving scenes.
IV-D Time-Dense Optical Flow Evaluation
@@
Figure 10: Illustration of our time-dense optical flow evaluation. Each row in the table shows the results of different time steps within a DSEC sample window of $100ms$ (total $14$ steps under $bins{=}15$). E-RAFT* indicates that the results at the intermediate time steps are obtained by interpolating the E-RAFT predictions over the complete time range.
As the DSEC flow dataset does not provide high frame-rate optical flow ground truth, we are not able to evaluate the precision of our model for temporally dense optical flow directly using the EPE metric. We propose a new unsupervised metric called RFWL to address this issue and assess the precision of event-based optical flow. Please refer to Sec. III-D for more details. Under the setting of $bins{=}15$, our model can obtain $14$ consecutive optical flow estimates for every sample from events occurring within $100ms$ on the DSEC test dataset. The output $V_{0,i}$ represents the optical flow estimation ranging from $0ms$ to $100{\times}\frac{i}{14}ms$. In contrast, E-RAFT is limited to obtaining single optical flow estimates within the time range of $0ms$ to $100ms$. Due to the absence of time-dense ground truth in the DSEC dataset, we directly compare our model with E-RAFT. We use time-dense optical flow results from our model and E-RAFT’s interpolated optical flow to generate motion-compensated frames for events during the relevant time periods. The demonstration of this process is shown in Fig. 10. We evaluate all test sequences for the DESC-Flow dataset and provide the detailed data in Tab. 3. Among the evaluated seven sequences, our model outperforms E-RAFT in terms of average RFWL in five sequences. Since movement in the DSEC dataset is predominantly linear, direct interpolation already yields relatively accurate intermediate optical flow. As a result, our method surpasses E-RAFT by small yet consistent margins on this dataset. Additionally, our runtime for a single optical flow estimation is only $10\%$ ($9.2ms$ vs. $93ms$) of E-RAFT’s runtime.
TABLE 4: Ablations of event representation ($bins{=}15$).

### v3

@@ 4.1 Datasets
DSEC. The DSEC dataset, introduced by Gehrig et al. Gehrig et al. (2021), consists of $24$ sequences captured in real-world outdoor driving scenes. The dataset includes a total of $7,800$ training samples and $2,100$ testing samples, captured at a resolution of $640{\times}480$. It covers both daytime and nighttime scenarios, encompassing small and large optical flows. Moreover, the DSEC dataset offers high-precision sparse ground truth values for optical flow. As the dataset does not include an official validation set, we divided the training dataset randomly using a fixed seed. To create the training and validation sets, we allocated them in a $4{:}1$ ratio. Consequently, only $80\%$ of the training data was utilized for training our model.
@@
MVSEC. MVSEC Zhu et al. (2018) is a classic real-world event optical flow dataset that comprises sequences from various indoor and outdoor scenes with a resolution of $346{\times}260$. The event density of the dataset is relatively sparse, and the optical flow ground truth distribution mostly highlights small flows. As a result, two different time intervals, i.e., $dt{=}1$ and $dt{=}4$, are used for accuracy evaluation. $dt{=}1$ means that adjacent grayscale frames are used as a sample, whereas $dt{=}4$ represents a four-fold increase in the grayscale frame interval. We follow the E-RAFT convention Gehrig et al. (2021), conduct training solely on the outdoor day 2 sequence, and evaluate on $800$ samples data from outdoor day 1.
@@ 4.2 Implementation Details
For the DSEC dataset Gehrig et al. (2021), we use Adam optimizer with a batch size of $6$ and a learning rate of $5{\times}1e^{-4}$ to train for $100k$ iterations on the training set we divided, which accounts for $80\%$ of the total dataset. Then, we reduce the learning rate by a factor of $10$ and continue training for another $10k$ iterations. During the model training on the DSEC dataset, two online data augmentation methods, namely random cropping and horizontal flipping, were employed. Random cropping was performed with a size of $288{\times}384$, while horizontal flipping was applied with a probability of $50\%$.
@@
For the MVSEC dataset Zhu et al. (2018), we employ the Adam optimizer with a batch size of $3$ and a learning rate of $1{\times}1e^{-3}$ to train for $120k$ iterations. We utilize random cropping with a size of $256{\times}256$ and horizontal flipping with a probability of $50\%$ during training. For MVSEC scenes with $dt{=}1$, we use the setting with #$\#bins{=}3$; for MVSEC with $dt{=}4$, we use the setting with #$\#bins{=}9$.
Table 1: Evaluation results on the DSEC-Flow’s test set Gehrig et al. (2021).
{adjustwidth}
[…]
4.3 Regular Flow Evaluation Prototype
Evaluation on DSEC. Tab. 1 presents a comprehensive evaluation on the DSEC-Flow benchmark Gehrig et al. (2021), including performance and computational complexity metrics. The metrics assess accuracy via EPE (L2 endpoint error in pixels), AE (angular error in degrees), and nPE (percentage of pixels with optical flow magnitude error greater than n pixels); model complexity via Params (millions of parameters); and computational cost via GMACs (Giga Multiply-Accumulate operations per inference, in billions).
Analyzing the results, our EVA-Flow establishes itself as a leading solution for time-dense optical flow. Compared to other approaches capable of high-frequency output (Taming_CM, LSTM-FlowNet), EVA-Flow achieves superior temporal characteristics with the lowest input latency (5ms) and the highest prediction rate (200Hz). While delivering this enhanced temporal resolution, it maintains a competitive level of accuracy. Although slightly less accurate than the top-performing non-time-dense methods, EVA-Flow demonstrates an excellent trade-off between high-frequency prediction capabilities and flow estimation precision. Crucially, EVA-Flow achieves this performance with remarkable computational efficiency, making it highly suitable for resource-constrained applications.
Importantly, the reported GMACs reflect the cost for generating a single optical flow prediction. This allows a fair comparison of per-prediction efficiency, irrespective of the output rate (e.g., our 200Hz vs. 10/100Hz for others). EVA-Flow’s low per-prediction cost is what enables its high-frequency operation.
E-RAFT Gehrig et al. (2021), a representative method for event-based optical flow, incorporates the RAFT’s Teed and Deng (2020) superior design features, such as correlation computation and iterative optical flow estimation, resulting in enhanced optical flow accuracy. It serves as the foundation for the latest state-of-the-art methods, TMA Liu et al. (2023) and BFlow Gehrig et al. (2024). Consequently, we have chosen E-RAFT as a benchmark for our qualitative experiments.
@@ 4.3 Regular Flow Evaluation Prototype
The qualitative results of our method on the DSEC flow dataset are shown in Fig. 5. To make a more intuitive comparison of the accuracy of optical flow estimation, we also visualized the IWEs. It can be seen from the Regions of Interest (ROIs) in Fig. 5 that, compared with E-RAFT, our model can better distinguish the outline of small objects (such as poles) and obtain more accurate optical flow estimates. Our analysis suggests that the distinguishing feature of our approach, in contrast to E-RAFT, lies in the implicit use of warp-aligned event edges to extract motion features, whereas E-RAFT relies on a correlation volume. For small objects with simple, well-defined contours, warping and aligning event features is relatively straightforward. However, due to limited textures, the correlation-based method, which depends on measuring feature similarity, struggles to capture high-quality motion features in areas with insufficient texture, leading to the inferior performance of E-RAFT in such cases.
Table 2: Evaluation results on MVSEC Zhu et al. (2018).
$dt{=}1$ $dt{=}4$
EPE $\downarrow$ Outlier%$\downarrow$ EPE $\downarrow$ Outlier%$\downarrow$
SSL EV-FlowNet Zhu et al. (2018) 0.49 0.20 1.23 7.30
Spike-FlowNet Lee et al. (2020) 0.49 - 1.09 -
STE-FlowNet Ding et al. (2022) 0.42 0.00 0.99 3.90
Taming_CM Paredes-Vallés et al. (2023) 0.27 0.05 - -
USL Hagenaars et al. Hagenaars et al. (2021) 0.47 0.25 1.69 12.5
Zhu et al. Zhu et al. (2019) 0.32 0.00 1.30 9.70
Shiba et al. Shiba et al. (2022) 0.30 0.10 1.25 9.21
SL TMA Liu et al. (2023) 0.25 0.07 0.7 1.08
E-RAFT Gehrig et al. (2021) 0.24 0.00 0.72 1.12
EVA-Flow (Ours) 0.25 0.00 0.82 2.41
E-RAFT Gehrig et al. (2021) (Zero-Shot) 0.53 1.42 1.93 17.7
EVA-Flow (Zero-Shot) 0.39 0.07 0.96 4.92
@@
Figure 5: Qualitative examples of optical flow predictions on the DSEC-Flow test sequences. The four samples are taken from distinct scenes within different sequences. Regions of interest, marked with red boxes, are magnified 2×2$\times$2 × and shown in the top-left corner of each image.
@@
Figure 6: Qualitative examples of optical flow predictions on the outdoor day 1 sequence of the MVSEC dataset Zhu et al. (2018) ($dt{=}4$). The model used here is only trained on the DSEC dataset Gehrig et al. (2021). The respective EPE and RFWL metrics can be found in the bottom left corner of the images. The areas of interest are denoted as red boxes in the images.
@@
Evaluation on the MVSEC dataset. The evaluation results of the MVSEC benchmark Zhu et al. (2018) are presented in Tab. 2. SSL, USL, and SL stand for self-supervised-, unsupervised-, and supervised learning, respectively. While our model’s accuracy is slightly lower than that of E-RAFT based on the data in the table, it is worth highlighting that our zero-shot results display a significantly better accuracy when directly testing the performance of the DSEC-trained model on the MVSEC dataset. Especially noteworthy is our model’s performance at $dt{=}4$, where our model’s EPE is halved compared to that of E-RAFT ($0.96$ vs. $1.93$). These results demonstrate the superior generalizability of the proposed EVA-Flow.
@@
Figure 7: Qualitative examples of optical flow predictions on our EVA-FlowSet. The model used here is only trained on the DSEC dataset Gehrig et al. (2021). The initial two rows of samples are obtained from high-speed sequences, while the remaining two rows are taken from normal-speed ones. Best view via zoom in.
@@
The qualitative results of testing MVSEC under the zero-shot setting, i.e., the model is only trained on DSEC, can be found in Fig. 6. Owing to the small optical flow of MVSEC, the motion compensation effect is indistinct under the $dt{=}1$ setting. Hence, we apply a longer time range of grayscale frames, i.e., $dt{=}4$. There is a marked discrepancy in motion speed between the MVSEC and DSEC datasets, as depicted in Fig. 6 and Fig. 5. Nonetheless, our model is capable of achieving near-ground-truth optical flow results on the MVSEC dataset, while E-RAFT shows inadequate generalization performances, particularly in non-event ground areas. Additionally, our model demonstrates superior capabilities for small targets such as poles and road markings, accurately depicting their contours with improved motion compensation results.
@@
Evaluation on EVA-FlowSet. Qualitative results of the curve motion scene dataset EVA-FlowSet can be found in Fig. 7. By visualizing the optical flow images, it can be observed that, for both fast-moving scenes and normal scenes, our model can estimate the contour of moving objects more accurately than E-RAFT. The magnified ROI area shows that our model obtains sharper motion compensation results in the first three rows of fast-moving scenes, while E-RAFT’s motion compensation performance is comparable to our model in the last three rows of moderate-moving-speed scenes. E-RAFT calculates the correlation between the event voxel grid before and after the start time to estimate the optical flow, which implicitly assumes that the optical flow is constant throughout the entire time. In contrast, our method uses the SMR module iteratively in each time step to obtain continuous optical flow estimation, which is suitable for curve motion scenes. In scenes with moderate movement speeds, even if the motion trajectory is curved, the motion can be approximated as linear due to the relatively slow speed. Therefore, in cases of moderate movement speeds, the motion compensation results of E-RAFT are comparable to those of our model. However, in fast-moving scenes, the linear motion assumption of E-RAFT becomes invalid due to the inherently curved nature of the motion trajectory. In contrast, our model, which employs a dense optical flow estimation framework independent of the linear motion assumption, significantly outperforms E-RAFT in fast-moving scenarios.
4.4 Time-Dense Optical Flow Evaluation
@@
Figure 8: Illustration of time-dense optical flow evaluation. Each row presents results from different time steps within a $100\,ms$ DSEC sample window (14 steps, #$\#bins{=}15$). E-RAFT* denotes interpolated predictions across intermediate time steps.
Table 3: Comparison of anytime flow estimation on DSEC-Flow.
{adjustwidth}
[…]
With #$\#bins{=}15$, our model produces $14$ consecutive optical flow estimates from events within a $100\,ms$ window on the DSEC test set. The output $F_{0,i}$ covers $0\,ms$ to $100{\times}\frac{i}{14}\,ms$, while E-RAFT provides only a single estimate for the entire $0\,ms$ to $100\,ms$ range. Due to the absence of time-dense ground truth in the DSEC dataset, a direct comparison with E-RAFT is not feasible. To address this, we generate Images of Warped Events (IWEs) using our model’s time-dense flow estimates and E-RAFT’s interpolated flow, and evaluate them with RFWL (refer to Sec. 3.4). This process is illustrated in Fig. 8. All DSEC test sequences are evaluated, and detailed results are presented in Tab. 3.
Since $F_{0,i}$ represents forward optical flow from $0\,ms$ to $100{\times}\frac{i}{14}\,ms$, we use this to warp the events within that period to generate the corresponding IWE. Our model’s optical flow estimates at intermediate time steps are generally more accurate than E-RAFT’s interpolated results, leading to better RFWL performance. However, as $i$ approaches the final time step (e.g., $i{=}14$), E-RAFT’s more accurate final optical flow results in a higher RFWL. Nevertheless, across seven sequences, our method surpasses E-RAFT in average RFWL on five sequences, demonstrating that our approach reliably estimates high-frame-rate optical flow even without supervision from high-frame-rate ground truth.
{adjustwidth}
[…]
Figure 9: Evaluation results of time-dense optical flow using the RFWL in EVA-FlowSet. Models utilized here are exclusively trained on the DSEC dataset Gehrig et al. (2021).
We conducted a comparative analysis using the EVA-FlowSet, which comprises sequences featuring curved movements. The results, depicted in Fig. 9, indicate a significant performance improvement of our approach compared to E-RAFT. Since the EVA-FlowSet lacks ground truth optical flow, both E-RAFT and our model are only trained on the DSEC dataset. Fig. 9 presents a visual representation of the evaluation results comparing our method with E-RAFT in real-world scenarios with varying motion speeds. Specifically, in two fast-moving sequences, our model demonstrates a notably superior intermediate RFWL compared to E-RAFT. We observe that the overall accuracy of EVA-Flow is significantly higher during periods of rapid motion, yet it experiences a slight decrease in accuracy towards the end of the event bin cycle. This phenomenon correlates with the aforementioned results of the DSEC dense-time validation experiments. In the other two scenarios characterized by normal-speed sequences, our model’s RFWL performance is comparable to that of E-RAFT. This equivalence arises because, at insufficient movement speeds along a curved path, the motion can be briefly approximated as linear.
We also assessed the efficacy of our dense temporal optical flow methodology from an alternative perspective. In the DSEC-Flow, where the ground truth flow frame rate is 10 Hz and each sample spans $100\,ms$, we created a $200\,ms$ ground truth for event optical flow by warping together the preceding and subsequent $100\,ms$ optical flows of our validation set. Thereafter, we applied a model trained on the original dataset, which consists of $100\,ms$ samples, to estimate this extended validation set. Tab. 4 (right side) displays the outcomes of this experiment. For the EVA-Flow model, which was trained with Bins set to $b$, we used the adjusted setting of $b^{\prime}=1+(b-1)\times 2$ for estimating the $200\,ms$ optical flow. This approach maintains the duration of each bin as it was during training and aligns the bin lengths with the extended overall estimation period. Conversely, E-RAFT, when processing a $200\,ms$ sample, adheres to its original configuration of $15$ bins. The data in Tab. 4 show that our model’s error did not escalate when estimating the validation set with an extended time span; rather, its precision improved, with the average End Point Error (EPE) being less than double that of the test set with $100\,ms$ samples. In comparison, E-RAFT’s EPE surpassed three times the EPE of the $100\,ms$ sample test set.
We also observed similar results on the MVSEC dataset. Tab. 2 displays the zero-shot results for E-RAFT and our model, which was trained on DSEC and evaluated on MVSEC with #$\#bins{=}15$. When assessing optical flow estimation on the MVSEC dataset with our model, we use the identical checkpoint but adjust the bin count according to varying $dt$ values. Due to the difference in motion speeds between the slow-motion MVSEC and the fast-motion DSEC datasets, we set #$\#bins{=}4$ for $dt{=}1$ and #$\#bins{=}13$ for $dt{=}4$. It is apparent that our model achieves a lower endpoint error (EPE) of $0.96$ at $dt{=}4$, which is less than four times the EPE at $dt{=}1$ ($0.39$).
Interestingly, the results indicate that as we progressively iterate and generate optical flow for longer time intervals, the average per-unit-time optical flow error does not increase, but rather decreases. These findings provide clear evidence for the efficacy of our framework in estimating time-dense optical flow. Furthermore, it underscores a significant advantage of the framework’s adaptable nature: the flexibility to adjust the number of prediction bins. This adjustment minimizes the difference in event rates per bin between the training and prediction phases, thereby enhancing accuracy during deployment.
Table 4: Experiments of different #bins on DSEC-Flow.
Test Sequences of $10Hz$ Validation Split of $5Hz$
Model #Bins (Training) EPE $\downarrow$ AE $\downarrow$ 1PE $\downarrow$ 3PE $\downarrow$ #Bins (Evaluating) EPE $\downarrow$ 1PE $\downarrow$ 3PE $\downarrow$ 5PE $\downarrow$
E-RAFT 15 0.79 2.85 12.7 2.7 15 2.96 43.5 17.6 10.3
6 0.955 3.29 16.7 3.9 11 1.73 42.9 14.7 7.9
11 0.926 3.34 16.1 3.5 21 1.86 51.1 15.8 8.0
15 0.895 3.39 16.1 3.3 29 1.82 49.4 15.6 8.0
21 0.877 3.31 15.9 3.2 41 1.89 48.7 16.8 8.8
EVA-Flow 31 0.901 3.37 17.0 3.2 61 2.02 53.5 18.2 9.4
4.5 Ablation Study
Figure 10: Impact of inference number of UVG bins (N) on Time-Dense Flow Quality (EVA-FlowSet, $T=20ms$).
Unified Voxel Grid. Tab. 5 compares the results of our framework on the DSEC test set using Voxel Grid Zhu et al. (2019) and our proposed Unified Voxel Grid representation. The EPE of our proposed Unified Voxel Grid is $7.3\%$ lower than that of Voxel Grid. This difference stems from the fact that each event bin in the Unified Voxel Grid adheres to a consistent event representation, while the representation patterns of the first and last bins in the Voxel Grid do not match those of the middle bins (refer to Fig. 3).
Table 5: Ablations of event representation (#$\#bins{=}15$).